OpenStudy (anonymous):

Ok, got another "monster" LOL! Anyone want to help me with this one: A = P(1 + r) t (where t is an exponent. t =time, P = principal invested and r is rate in decimal. t =2, P=$300 and r=9% or .09? Much appreciated, denise

OpenStudy (anonymous):

A= 300(1+.09)^2 = 300(1.09)^2 =300*1.1881 $356.43 did i miss the monster part?

OpenStudy (mathstudent55):